laura has $4.50 in dimes and quarters. She has three more dimes than quarters. How many quarters does she have?
Answer by checkley79(3341)

d=q+3
.25q+.10d=2.40
.25q+.10(q+3)=2.40
.25q+.10q+.30=2.40
.35q=2.40-.30
.35q=2.10
q=2.10/.35
q=6 quarters.
6+3=9 dimes.
Proof:
.25*6+.10*9=2.40
1.50+.90=2.40
2.40=2.40